# bibliography.bib

### v1

@@ Abstract
Prior-data fitted networks (PFNs) have recently been proposed as a promising way to train tabular foundation models. PFNs are transformers that are pre-trained on synthetic data generated from a prespecified prior distribution and that enable Bayesian inference through in-context learning. In this paper, we introduce CausalFM, a comprehensive framework for training PFN-based foundation models in various causal inference settings. First, we formalize the construction of Bayesian priors for causal inference based on structural causal models (SCMs) in a principled way and derive necessary criteria for the validity of such priors. Building on this, we propose a novel family of prior distributions using causality-inspired Bayesian neural networks that enable CausalFM to perform Bayesian causal inference in various settings, including back-door, front-door, and instrumental variable adjustment. Finally, we instantiate CausalFM and explicitly train a foundation model for estimating conditional average treatment effects (CATEs) using back-door adjustment. We show that CausalFM performs competitively for CATE estimation using various synthetic and semi-synthetic benchmarks. In sum, our framework can be used as a general recipe to train foundation models for various causal inference settings. In contrast to the current state-of-the-art in causal inference, CausalFM offers a novel paradigm with the potential to fundamentally change how practitioners perform causal inference in medicine, economics, and other disciplines.
@@ 1 Introduction
Causal inference is a cornerstone of empirical research in disciplines such as economics [Angrist.1990, Imbens.1994], medicine [Feuerriegel.2024, Glass.2013], and marketing [Varian.2016]. It enables the estimation of causal effects from observational and randomized data, which is essential for reliable decision-making. In personalized medicine, for instance, it supports identifying the most effective treatment by predicting patient outcomes under different therapeutic options.
@@
In recent years, machine learning, and especially deep learning methods, have gained significant traction in causal inference (e.g., [Curth.2021, Melnychuk.2022, Schweisthal.2023, Shalit.2017, Shi.2019]). These methods offer several advantages for causal effect estimation in practice, including the ability to handle large-scale and high-dimensional datasets with complex confounding structures and to model heterogeneity of causal effects [Feuerriegel.2024]. However, most existing approaches require retraining a model for each new causal task or dataset. To this end, existing approaches lack the flexibility to perform test-time inference without additional retraining, which limits their practicality in real-world settings.
@@
Meanwhile, foundation models have emerged as a transformative paradigm in machine learning [devlin2018bert, touvron2023llama, touvron2023llama2, lahat2024assessing], which offer a key advantage in that they allow for flexible, test-time inference without retraining. These models are pre-trained on large datasets and can generalize across tasks and domains. Examples include large language models (LLMs) in natural language processing and vision transformers in computer vision. However, this paradigm shift toward test-time inference has not yet had a comparable impact on causal inference. Most current approaches in causal machine learning still rely on specialized models tailored to specific tasks, requiring practitioners to manually select, train, and validate appropriate estimation methods for each new dataset.
@@
In this paper, we propose a change to the paradigm for causal inference based on the idea of foundation models trained for tabular causal inference. For this, we build on the recently proposed prior-data fitted networks (PFNs) [Muller.2022, Hollmann.2023], which are transformers pre-trained on purely synthetic datasets generated from a prespecified prior distribution. PFNs enable Bayesian inference purely through in-context learning, allowing for flexible and efficient predictions without requiring additional training at inference time [Nagler.2023]. While recent works have demonstrated the effectiveness of tabular foundation models based on PFNs for various tasks, only two concurrent works have proposed PFNs tailored for causal inference [Balazadeh.2025, Robertson.2025]. However, these are either restricted to specific causal inference settings (namely, only back-door adjustment) or do not offer identifiability guarantees.
@@
We introduce CausalFM, a comprehensive framework for training PFN-based foundation models for various causal inference settings. For this purpose, we introduce the CausalFM prior: a novel family of prior distributions based on structural causal models respecting the underlying causal inference problem at hand. We first formalize and derive necessary criteria on how to construct such SCM-based priors for causal inference in principle. Then, we propose a concrete instantiation using Bayesian neural networks and provide a learning algorithm that leverages the SCM’s ability to simulate counterfactual data and perform consistent Bayesian inference in a wide range of causal inference settings.
@@
Compared to classical causal inference methods, foundation models trained based on our CausalFM offer the following advantages: (i) There is no need for additional training at inference time as our CausalFM performs inference entirely through in-context learning, enabling fast and flexible deployment across new datasets. (ii) The Bayesian nature of our CausalFM provides principled uncertainty quantification, which is critical for downstream decision-making and for detecting situations with poor treatment overlap. (iii) The model automatically learns to select among alternative adjustment strategies such as back-door, front-door, or instrumental variable (IV) adjustment based on the data distribution and task at hand. For instance, in the presence of a weak instrument, the model may automatically choose to avoid IV adjustment in favor of a more robust alternative, thereby improving estimation performance and reducing variance. (iv) Our CausalFM builds upon rigorous identifiability guarantees to ensure asympotically valid inferences.
@@
Our contributions [^1] Code is available at https://github.com/yccm/CausalFM. are: (1) We formalize the constructions of priors based on structural causal models (SCMs) for Bayesian causal inference and derive necessary conditions for their validity. (2) We propose an explicit CausalFM prior based on Bayesian neural networks that are compatible with the structure of the causal inference problem at hand. We also propose a learning algorithm to train PFNs for causal inference problems that leverages our CausalFM prior to simulate counterfactuals to mitigate the fundamental problem of causal inference. (3) We propose a concrete instantiation of our framework: we train a PFN for estimating conditional average treatment effects (CATEs) via back-door adjustment. We show empirically that CausalFM performs competitively and outperforms current state-of-the-art CATE estimators on a variety of benchmarks.
@@ 2 Related Work
Prior-data-fitted networks (PFNs) as tabular foundation models.
Foundation models are an emerging paradigm that has revolutionized machine learning for various data modalities, particularly language and vision tasks [devlin2018bert, touvron2023llama, touvron2023llama2, lahat2024assessing]. The same paradigm is now being explored for tabular data—a modality that underpins the large majority of analyses in science and business [vanBreugel.2024].
A powerful approach is given by prior-data-fitted networks (PFNs) [Muller.2022], which are large transformers trained exclusively on synthetic data sampled from a user-specified prior distribution. Then, at test time, it performs Bayesian inference through in-context learning by approximating the posterior predictive distribution. TabPFN [Hollmann.2023, Hollmann.2025] scaled this idea by pairing the transformer with a Bayesian neural-network prior over structural causal models (SCMs) and demonstrating state-of-the-art performance on various tabular benchmarks. Subsequent work extended PFNs to time–series forecasting [Hoo.2025] and analyzed their in-context learning abilities theoretically [Nagler.2023]. Critically, all existing PFNs are trained only for predictive tasks and do not target causal estimands; they therefore are not designed for causal inference of treatment effects, which is the goal of our paper.
Causal inference with machine learning.
Causal inference, such as the estimation of average treatment effects, originates from fields like econometrics [Imbens.1994, Angrist.1990], statistics [vanderLaan.2006], and epidemiology [Robins.1986, Robins.1994]. More recently, machine learning has focused on estimating heterogeneous effects to support personalized decision-making.
One line of work are frequentist methods, which build on semiparametric theory and efficient influence functions [Robins.1994b, Robins.1999], yielding model-agnostic estimators that are doubly robust and Neyman-orthogonal [vanderLaan.2006b, Chernozhukov.2018, Nie.2021, Foster.2023, Kennedy.2023b]. Another line of work builds upon specific machine learning methods/ architectures such as regression trees [Wager.2018] or neural networks [Johansson.2016, Shalit.2017, Shi.2019] and adopts them to causal inference. Bayesian alternatives include Bayesian additive regression trees [Hahn.2020] or Gaussian-process counterfactual regression [Alaa.2017]. However, all of the existing estimators above must be retrained for every new dataset and causal setting. In contrast, our CausalFM allows to pre-train models to approximate Bayesian causal inference via in-context learning.
Amortized causal inference.
@@
Several recent papers pre-train large neural networks on synthetic data so that they can solve causal tasks via in-context learning during inference time. Examples include causal discovery [Mahajan.2025], ATE estimation under unconfoundedness [Zhang.2024b], zero-shot- and few-shot learning [Nilforoshan.2023, Iwata.2023], and reinforcement-learning [Lee.2023]. These methods validate the feasibility of treating causal inference as an in-context learning problem but remain restricted to specific causal inference settings – which typically do not allow to accommodate unobserved confounding.
@@
Black-box causal inference (BBCI) [Bynum.2025] takes a step towards pertaining models on synthetic data to perform causal inference in a variety of settings. However, their approach is different as compared to our CausalFM as follows: (i) BBCI does not build upon a Bayesian framework. In contrast, building upon PFNs allows us to perform approximate Bayesian causal inference and thus provide rigorous uncertainty quantification for treatment effects. (ii) The proposed data-generating processes in BBCI are not tailored for high-dimensional causal inference settings (as the authors mention in their Sec. 7). In contrast, our CausalFM prior leverages Bayesian neural networks inspired by TabPFN [Hollmann.2023] to create high-dimensional SCM-based prior distributions. (iii) Beyond proposing a new method, we provide novel formalizations and theoretical results of constructing valid SCM-based priors for Bayesian causal inference.
@@
Foundation models for causal inference: We are aware of only two works published concurrently to this paper that propose PFN-based models for causal inference, but each with clear limitations: (i) [Balazadeh.2025] proposes a PFN similar to ours, but it is restricted to only back-door adjustment, i.e., imposes the unconfoundedness assumption throughout their paper. In contrast, we propose a framework for constructing PFN-based foundation models for a large class of causal inference problems, including both front-door adjustment and instrumental variable settings with unobserved confounding. As a consequence, our Bayesian prior construction for training the PFN is different and allows to sample various causal inference settings. (ii) [Robertson.2025] proposes to train a single PFN on various different causal inference settings without providing identifiability assumptions to the model. We will show later that the approach of [Robertson.2025] has a crucial drawback: because the causal quantity of interest is not identified, the PFN learns a posterior that may never concentrate around the true causal quantity, thus leading to asymptotically biased estimators. In contrast, we propose to train separate PFNs for different identifiable causal inference settings. As such, we follow established philosophy in causal inference that separates identifiability and estimation steps [Pearl.2009]: the identifiability step should be established by the practitioner using domain knowledge (e.g., establishing whether a certain variable is a valid instrument), while the estimation step can be treated as a purely statistical learning problem and thus can be handed over to a pre-trained foundation model.
@@ 3.1 Background on PFNs
In tabular prediction problems, one considers a population $(X,Y)\sim\mathbb{P}\in\mathcal{P}$. Given a finite sample $\mathcal{D}_{n}\sim\mathbb{P}$ of size $n$, the goal is to estimate the conditional distribution $\mathbb{P}(Y=y\mid X=x)$. PFNs formulate this task in a Bayesian non-parametric way by placing a prior distribution $\Pi$ on $\mathcal{P}$, i.e., a prior over data-generating distributions [Muller.2022, Nagler.2023]. Sampling proceeds hierarchically via
$\mathbb{P}\sim\Pi,\qquad(X_{i},Y_{i})\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathbb{P}.$ (1)
Then, Bayes’ rule yields the posterior distribution
$\Pi(\mathbb{P}\mid\mathcal{D}_{n})\;\propto\;\Pi(\mathcal{D}_{n}\mid\mathbb{P})\,\Pi(\mathbb{P}),$ (2)
where $\Pi(\mathcal{D}_{n}\mid\mathbb{P})$ is the likelihood of the sample $\mathcal{D}_{n}$ under $\mathbb{P}$ and $\propto$ denotes proportionality up to a multiplicative constant. The corresponding posterior-predictive distribution is defined as
$\Pi(Y\mid\mathcal{D}_{n},x)=\int\mathbb{P}(Y\mid X=x)\,\Pi(\mathbb{P}\mid\mathcal{D}_{n})\,\mathop{}\!\mathrm{d}\mathbb{P},$ (3)
which is the probability of $Y$ given test point $x$ and observed data $\mathcal{D}_{n}$.
@@
PFNs are neural networks $q_{\theta}(Y\mid\mathcal{D}_{n},x)$ that parameterize the family of predictive posterior distributions with trainable parameters $\theta$. That is, PFNs map the entire dataset $\mathcal{D}_{n}$ and a query $x$ to a distribution over $\mathcal{Y}$. In terms of architecture, PFNs are permutation-equivariant transformers [Vaswani.2017] as they allow for scalable training and leverage the attention mechanism to effectively extract information from $\mathcal{D}_{n}$.
PFNs are trained by minimizing the negative log-likelihood loss
$\mathcal{L}(\theta)=\mathbb{E}_{N\sim\Pi_{N}}\!\bigl{[}\mathbb{E}_{\mathbb{P}\sim\Pi}\bigl{[}-\log q_{\theta}(Y\mid X,\mathcal{D}_{N})\bigr{]}\bigr{]},$ (4)
where $\Pi_{N}$ is a prior on the sample sizes. In practice, we sample a sample size $N_{j}\sim\Pi_{N}$, a probability distribution $\mathbb{P}_{j}\sim\Pi$, a dataset $\mathcal{D}_{N_{j}}^{j}\sim\mathbb{P}_{j}$, and test points $(x_{j},y_{j})\sim\mathbb{P}_{j}$ and then approximate the PFN loss via
^$\hat{\mathcal{L}}(\theta)=\sum_{j}\bigl{[}-\log q_{\theta}(y_{j}\mid\mathcal{D}_{N_{j}}^{j},x_{j})\bigr{]},$ (5)
@@
which is consistent for the exact posterior-predictive under regularity conditions [Nagler.2023]. Note that all training data are synthetic, i.e., sampled from the prior $\Pi$. Furthermore, the trained PFN can be deployed on arbitrary real datasets without further training.
3.2 Task

### v2

# Foundation Models for Causal Inference via Prior-Data Fitted Networks

@@ Abstract
Prior-data fitted networks (PFNs) have recently been proposed as a promising way to train tabular foundation models. PFNs are transformers that are pre-trained on synthetic data generated from a prespecified prior distribution and that enable Bayesian inference through in-context learning. In this paper, we introduce CausalFM, a comprehensive framework for training PFN-based foundation models in various causal inference settings. First, we formalize the construction of Bayesian priors for causal inference based on structural causal models (SCMs) in a principled way and derive necessary criteria for the validity of such priors. Building on this, we propose a novel family of prior distributions using causality-inspired Bayesian neural networks that enable CausalFM to perform Bayesian causal inference in various settings, including for back-door, front-door, and instrumental variable adjustment. Finally, we instantiate CausalFM and explicitly train models to perform in-context learning in these settings. We show that CausalFM achieves competitive in-context learning performance even when compared to baselines that are specifically trained for the task at hand. In sum, our framework can be used as a general recipe to train foundation models for various causal inference settings. In contrast to the current state-of-the-art in causal inference, CausalFM offers a novel paradigm with the potential to fundamentally change how practitioners perform causal inference in medicine, economics, and other disciplines.
@@ 1 Introduction
Causal inference is a cornerstone of empirical research in disciplines such as economics (Angrist, 1990; Imbens & Angrist, 1994), medicine (Feuerriegel et al., 2024; Weberpals et al., 2025), and marketing (Varian, 2016). It enables the estimation of causal effects from observational and randomized data, which is essential for reliable decision-making (Kern et al., 2025). In personalized medicine, for instance, it supports identifying the most effective treatment by predicting patient outcomes under different therapeutic options.
@@
In recent years, machine learning, and especially deep learning methods, have gained significant traction in causal inference (Curth & van der Schaar, 2021; Ma et al., 2025; 2024; Melnychuk et al., 2022; Schweisthal et al., 2023; Shalit et al., 2017a; Shi et al., 2019)). These methods offer several advantages for causal effect estimation in practice, including the ability to handle large-scale and high-dimensional datasets with complex confounding structures and to model heterogeneity of causal effects (Feuerriegel et al., 2025). However, most existing approaches require retraining a model for each new dataset. To this end, existing approaches lack the flexibility to perform inference for new datasets without additional retraining, which limits their practicality in real-world settings.
@@
Meanwhile, foundation models have emerged as a transformative paradigm in machine learning (Devlin, 2018; Lahat et al., 2024; Touvron et al., 2023b; a), which offer a key advantage in that they allow for flexible, test-time inference without retraining. These models are pre-trained on large datasets and can generalize across tasks and domains. Examples include large language models (LLMs) in natural language processing and vision transformers in computer vision. However, this paradigm shift toward test-time inference has not yet had a comparable impact on causal inference. Most current approaches in causal machine learning still rely on specialized models tailored to specific tasks, requiring practitioners to manually select, train, and validate appropriate estimation methods for each new dataset.
@@
In this paper, we propose a change to the paradigm for causal inference based on the idea of foundation models trained for tabular causal inference. For this, we build on the recently proposed prior-data fitted networks (PFNs) (Müller et al., 2022; Hollmann et al., 2023), which are transformers pre-trained on purely synthetic datasets generated from a prespecified prior distribution. PFNs enable Bayesian inference purely through in-context learning, allowing for flexible and efficient predictions without requiring additional training for new tasks (Nagler, 2023). While recent works have demonstrated the effectiveness of tabular foundation models based on PFNs for various tasks, only two concurrent works have proposed PFNs tailored for causal inference (Balazadeh et al., 2025; Robertson et al., 2025). However, these are either restricted to specific causal inference settings (namely, only back-door adjustment) or do not offer identifiability guarantees.
@@
We introduce CausalFM, a comprehensive framework for training PFN-based foundation models for various causal inference settings. For this purpose, we introduce CausalFM priors: a novel family of prior distributions based on structural causal models respecting the underlying causal inference problem at hand. We first formalize and derive necessary criteria on how to construct such SCM-based priors for causal inference in principle. Then, we propose a concrete instantiation using Bayesian neural networks and provide a learning algorithm that leverages the SCM’s ability to simulate interventional data to perform Bayesian inference in various causal inference settings.
@@
Compared to classical causal inference methods, models trained based on our CausalFM offer the following advantages: (i) There is no need for additional training for new datasets as our CausalFM performs inference entirely through in-context learning, enabling fast and flexible deployment across new datasets. (ii) The Bayesian nature of our CausalFM provides principled uncertainty quantification, which is critical for downstream decision-making and for detecting situations with poor treatment overlap. (iii) The model automatically learns to “select” an identifiability formula based on the data distribution and task at hand. (iv) Our CausalFM builds upon rigorous identifiability guarantees to ensure valid causal inference.
@@
Our contributions [^1] Code is available at https://github.com/yccm/CausalFM. are: (1) We formalize the constructions of priors based on structural causal models (SCMs) for Bayesian causal inference and derive necessary conditions for their validity. (2) We propose an explicit CausalFM prior based on Bayesian neural networks that are compatible with the structure of the causal inference problem at hand. We also propose a learning algorithm to train PFNs for causal inference problems that leverages our CausalFM prior to simulate counterfactuals to mitigate the fundamental problem of causal inference. (3) We propose concrete instantiations of our framework by training PFNs for estimating conditional average treatment effects (CATEs) in different causal inference settings. We show empirically that CausalFM performs competitively and outperforms current state-of-the-art CATE estimators on a variety of benchmarks.
@@ 2 Related Work
We provide an overview of related literature streams. Additional related work is in Appendix A.
@@
Amortized causal inference. Several recent papers pre-train large neural networks on synthetic data so that they can solve causal tasks via in-context learning. Examples include causal discovery (Mahajan et al., 2025), ATE estimation under unconfoundedness (Zhang et al., 2024), zero-shot- and few-shot learning (Nilforoshan et al., 2023; Iwata & Chikahara, 2023), and reinforcement-learning (Lee et al., 2023). These methods validate the feasibility of treating causal inference as an in-context learning problem but remain restricted to specific causal inference settings, which typically do not allow accommodating unobserved confounding.
@@
Black-box causal inference (BBCI) (Bynum et al., 2025) proposes synthetically-pretrained models to perform causal inference in a variety of settings. However, their approach is different: (i) BBCI does not build upon a Bayesian framework. In contrast, building upon PFNs allows us to perform approximate Bayesian causal inference and thus provide rigorous uncertainty quantification. (ii) The proposed data-generating processes in BBCI are not tailored for high-dimensional causal inference settings (as the authors mention in their Sec. 7). In contrast, our CausalFM prior leverages Bayesian neural networks inspired by TabPFN (Hollmann et al., 2023) to create SCM-based prior distributions. (iii) Beyond proposing a new method, we provide novel formalizations and theoretical results of constructing valid SCM-based priors for Bayesian causal inference.
Table 1: Overview of identifiability of PFN-based frameworks for causal inference.
Framework Backdoor Frontdoor IV
CausalPFN Balazadeh et al. (2025)
Do-PFN Robertson et al. (2025)
Ours (CausalFM)
@@
PFNs for causal inference: We are aware of only two concurrent works that propose PFN-based models for causal inference, but each with clear limitations (see Figure 1): (i) (Balazadeh et al., 2025) proposes a PFN similar to ours, but it is restricted to only back-door adjustment, i.e., imposes the unconfoundedness assumption throughout their paper. In contrast, we propose a framework for constructing PFN-based foundation models for a large class of causal inference problems, including both front-door adjustment and instrumental variable settings with unobserved confounding. (ii) Robertson et al. (2025) proposes to train a single PFN on various different causal inference settings without providing identifiability assumptions to the model. We will show later that the approach of Robertson et al. (2025) has a crucial drawback: because the causal quantity of interest is not identified, the PFN learns a posterior that may never concentrate around the true causal quantity, thus leading to asymptotically non-informative estimators. In contrast, we propose to infuse our PFNs with identifiability assumptions required for informative causal inference. As such, we follow established philosophy in causal inference that separates identifiability and estimation steps (Kern et al., 2025; Pearl, 2009): the identifiability step should be established by the practitioner using domain knowledge (e.g., establishing whether a certain variable is a valid instrument), while the estimation step can be treated as a purely statistical learning problem.
@@ 3.1 Background on PFNs
In tabular prediction problems, one considers a population $(X,Y)\sim\mathbb{P}\in\mathcal{P}$. Given a finite sample $\mathcal{D}_{n}\sim\mathbb{P}$ of size $n$, the goal is to estimate the conditional distribution $\mathbb{P}(Y=y\mid X=x)$. PFNs formulate this task in a Bayesian non-parametric way by placing a prior distribution $\Pi$ on $\mathcal{P}$, i.e., a prior over data-generating distributions (Müller et al., 2022; Nagler, 2023). Sampling proceeds hierarchically via $\mathbb{P}\sim\Pi$ and i.i.d. data $(X_{i},Y_{i})\sim\mathbb{P}$. Then, Bayes’ rule yields the posterior distribution $\Pi(\mathbb{P}\mid\mathcal{D}_{n})\;\propto\;\Pi(\mathcal{D}_{n}\mid\mathbb{P})\,\Pi(\mathbb{P})$, where $\Pi(\mathcal{D}_{n}\mid\mathbb{P})$ is the likelihood of the sample $\mathcal{D}_{n}$ under $\mathbb{P}$ and $\propto$ denotes proportionality up to a multiplicative constant. The corresponding posterior-predictive distribution is the probability of $Y$ given test point $x$ and observed data $\mathcal{D}_{n}$, i.e.,
$\Pi(Y\mid\mathcal{D}_{n},x)=\int\mathbb{P}(Y\mid X=x)\,\Pi(\mathbb{P}\mid\mathcal{D}_{n})\,\mathop{}\!\mathrm{d}\mathbb{P}.$ (1)
@@
PFNs are neural networks $q_{\theta}(Y\mid\mathcal{D}_{n},x)$ that parameterize the family of predictive posterior distributions with trainable parameters $\theta$. That is, PFNs map the entire dataset $\mathcal{D}_{n}$ and a query $x$ to a distribution over $\mathcal{Y}$. In terms of architecture, PFNs are permutation-equivariant transformers (Ashish Vaswani et al., 2017) as they allow for scalable training and leverage the attention mechanism to effectively extract information from $\mathcal{D}_{n}$. PFNs are trained by minimizing the negative log-likelihood loss $\mathcal{L}(\theta)=\mathbb{E}_{N\sim\Pi_{N}}\!\bigl[\mathbb{E}_{\mathbb{P}\sim\Pi}\bigl[-\log q_{\theta}(Y\mid X,\mathcal{D}_{N})\bigr]\bigr]$, where $\Pi_{N}$ is a prior on the sample sizes. In practice, we sample a sample size $N_{j}\sim\Pi_{N}$, a probability distribution $\mathbb{P}_{j}\sim\Pi$, a dataset $\mathcal{D}_{N_{j}}^{j}\sim\mathbb{P}_{j}$, and test points $(x_{j},y_{j})\sim\mathbb{P}_{j}$ and then approximate the PFN loss via
$\hat{\mathcal{L}}(\theta)=\sum_{j}\bigl[-\log q_{\theta}(y_{j}\mid\mathcal{D}_{N_{j}}^{j},x_{j})\bigr],$ (2)
@@
which is consistent for the exact posterior-predictive under regularity conditions (Nagler, 2023). Note that all training data are synthetic, i.e., sampled from the prior $\Pi$. Furthermore, the trained PFN can be deployed on arbitrary real datasets without further training.
3.2 Task: causal inference